OpenStudy (anonymous):

determine the x-value of the solution to the system of equations x+y=-3 x=y+19 a) x=-11 b)x=8 c)x=3 d)x=-3

mathslover (mathslover):

x+y = -3 x = -3-y so we have x =y +19 -3-y = y+19 -3 - 19 = y+y -22 = 2y

mathslover (mathslover):

we get y = -11 but given x + y = -3 so x + (-11) = -3 x-11 = -3 x = -3+11

mathslover (mathslover):

got it @albertcornejo ?

OpenStudy (anonymous):

oh okay i got it thanks.

mathslover (mathslover):

so what do u get for x?

OpenStudy (anonymous):

its 8

mathslover (mathslover):

good work..